OpenStudy (anonymous):

Line H is represented by the following equation: 2x + 2y = 8 What is most likely the equation for line K so the set of equations has infinitely many solutions? x + 2y = 2 x + y = 4 2x + y = 6 x − y = 8

OpenStudy (anonymous):

if they all have infinite many solutions then wouldnt that make them all the same line? @lil_chukie

OpenStudy (anonymous):

so rewrite every option they have into slope-intercept form

OpenStudy (anonymous):

I really don't get this concept

OpenStudy (anonymous):

@Prizzi

OpenStudy (anonymous):

ok im back

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

im going to solve it, then lead you to the answer, it may take a min.

OpenStudy (anonymous):

Ok, there is an answer to the problem!

OpenStudy (anonymous):

Since 2x+2y=8 and x+y=4, what do you notice about the two equations?

OpenStudy (anonymous):

ok, so 2x+2y=8 and x+y=4 have the relationship of the line being the same. Divide all the terms on both sides of 2x+2y=8 and you will get x+y=4!